OpenStudy (anonymous):

Solve by the substition method 7x-8y=-90 6x+83=y

OpenStudy (saifoo.khan):

x = -14 y =-1

OpenStudy (anonymous):

-14, -1

OpenStudy (cathyangs):

7x-8(6x+83)=-90 7x-48x-664=-90 -41x=574 x=-14. 7(-14)-8y=-90 -98-8y=-90 -8y=8 y=-1